OpenStudy (anonymous):

write un slope-intercept form. 2x+4y=8

OpenStudy (anonymous):

Simplifying 2x + 4y = 8 Solving 2x + 4y = 8 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-4y' to each side of the equation. 2x + 4y + -4y = 8 + -4y Combine like terms: 4y + -4y = 0 2x + 0 = 8 + -4y 2x = 8 + -4y Divide each side by '2'. x = 4 + -2y Simplifying x = 4 + -2y

OpenStudy (anonymous):

I was told we were supposed to have it y= not x=... But is that does seem right

OpenStudy (jdoe0001):

hint: solve for "y"

OpenStudy (triciaal):

@mathwiz916 please read the question before answering.

OpenStudy (anonymous):

ooops sorry lol

OpenStudy (triciaal):

@tb1 slope intercept format y = m x + b

OpenStudy (triciaal):

whatever you do to one side of the equation also do to the other to keep the balance

OpenStudy (triciaal):

when solving for any variable in any equation, gather all the terms with that variable to one side and simplify then divide each side by the coefficient of the variable

OpenStudy (triciaal):



OpenStudy (anonymous):

 i'm not sure, its been a while, are u trying to get y by itself?